OpenStudy (anonymous):

If sin x = .42, what is cos x?

OpenStudy (anonymous):

\[\sin ^{2}x+\cos ^{2}x=1\]

OpenStudy (anonymous):

\[\cos ^{2}x=1-\sin ^{2}x\]

OpenStudy (anonymous):

\[cosx=\sqrt{1-\sin^2}x\]

OpenStudy (anonymous):

find the arc sin of x first or sin(-1)0.42

OpenStudy (anonymous):

you get x = 24.8349

OpenStudy (anonymous):

then find the cos of that so cos 24.8349

OpenStudy (anonymous):

= 0.9075